OpenStudy (anonymous):

Find all solutions to the equation in the interval [0, 2π). sin 2x - sin 4x = 0

OpenStudy (anonymous):

sin(2x) = sin(4x) sin(2x) = sin(2*2x) sin(2x) = 2sin(2x)cos(2x) 2sin(2x)cos(2x) - sin(2x) = 0 sin(2x) * (2cos(2x) - 1) = 0 sin(2x) = 0 or cos(2x) = 1/2 x = 0 + n•π/2, π/6 + n•π, 5π/6 + n•π x = 0, π/6, π/2, 5π/6, π, 7π/6, 3π/2, 11π/6

OpenStudy (anonymous):

@FutureMathProfessor Why do you always give the answer away? That completely defeats the point of this website.

OpenStudy (anonymous):

I showed him how to do it, I think you just need to cool it with the spam

OpenStudy (anonymous):

Why did you bring in cos in step 3?

OpenStudy (anonymous):

It's an identity

OpenStudy (anonymous):

\[\bf \sin(2x)=2\sin(x)\cos(x) \implies \sin(2*2x)=sin(4x)=2\sin(2*x)\cos(2*x)\]Now re-write the equation:\[\bf \sin(2x)-2\sin(2x)\cos(2x)=0 \implies \sin(2x)(1-2\cos(2x))=0\]Can you find the x-values from here? @KirbyLegs

OpenStudy (anonymous):

I can looking at a unit circle. I get the process all the way up to the last step. Why is it 1-2cos(2x)

OpenStudy (anonymous):

Oh because it's 2sin?

OpenStudy (anonymous):

That's another identity

OpenStudy (anonymous):

Ok thanks guys. I guess i just needed to pull out sin2x

OpenStudy (anonymous):

Medal? :D